OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve. sin(x+y)=6x−6y at the point (π,π).

OpenStudy (anonymous):

i would like it if you told me the steps and then check my steps after i do them. thanks

OpenStudy (ranga):

Use chain rule to find the derivative of the first term.

OpenStudy (anonymous):

so the inner function would be x+y and the outside would be sinx?

OpenStudy (ranga):

Derivative of sin( something ) = cos( something) times derivative of (something)

OpenStudy (anonymous):

would the derivative of x+y=0

OpenStudy (ranga):

derivative of x with respect to x is 1 derivative of y with respect to x is dy/dx. But it is easier in implicit differentiations to write dy/dx as y'

OpenStudy (anonymous):

ok

OpenStudy (ranga):

So differentiate each term in the original equation first.

OpenStudy (anonymous):

cos(x+y)(1+y')=5-6y'?

OpenStudy (ranga):

Yes but the 5 on the right should be 6.

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

now i have cos(x+y)+y'cos(x+y)=6(1-y')

OpenStudy (ranga):

{cos(x+y)} (1+y') = 6 - 6y' There are two terms with y'. Group them together and express it as y' = something

OpenStudy (anonymous):

ok so cos(x+y)-6=6y'-y'cos(x+y)

OpenStudy (anonymous):

sorry -6y

OpenStudy (ranga):

-6y'

OpenStudy (anonymous):

yeah sorry i cant type right now

OpenStudy (ranga):

factor y' out on the right and express it as y' = something

OpenStudy (anonymous):

ok so it would be y'=\[\cos (x+y)-6/6+\cos(x+y)\]

OpenStudy (anonymous):

would that be -1

OpenStudy (ranga):

yes there should be a -outside or you can switch the numerator to: 6 - cos(x + y)

OpenStudy (anonymous):

ok so would that equal -1

OpenStudy (ranga):

would what equal -1? Answer?

OpenStudy (anonymous):

umm yes?

OpenStudy (ranga):

You get y' = { 6 - cos(x + y) } / { 6 + cos(x + y) } y' is the slope of the tangent line. Evaluate the slope at x = pi and y = pi Once you know the slope and a point on the line (pi, pi) you can find the equation of the tangent line at (pi, pi).

OpenStudy (anonymous):

is the slope 5/7

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

so i do the point slope form with (pi,pi) as the point?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

ok i got y=(5/7)x+(2/7)pi

OpenStudy (ranga):

Yes. Does it agree with the answer. You may have to put in the value for pi.

OpenStudy (anonymous):

yup its right. thanks

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

@ranga another problem i had was 2x^2+2xy+y^2=5 at the point (1,1). i got y'=-2x/(y+1) is that right?

OpenStudy (ranga):

I am getting y' = -(2x + y) / (x + y)

OpenStudy (anonymous):

oh ok. how

OpenStudy (ranga):

2x^2+2xy+y^2=5 Differentiate each term: 4x + (2xy' + 2y) + 2yy' = 0 4x + 2y + 2xy' + 2yy' = 0 2(2x + y) + 2y'(x + y) = 0 (2x + y) + y'(x + y) = 0 y' = -(2x + y) / (x + y)

OpenStudy (anonymous):

ok i did the 2xy derivative wrong

OpenStudy (ranga):

You got to apply product rule.

OpenStudy (anonymous):

ok i got it right now thanks

OpenStudy (ranga):

Glad to help.

OpenStudy (anonymous):

what would the derivative of y^2-4 be? 2yy'?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

and x^2-5=2x?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

ok thanks

OpenStudy (ranga):

sure.

OpenStudy (anonymous):

would the derivative of 5+x^4y=4x^3y+x^4y'?

OpenStudy (anonymous):

and what about arctan(xy)=1/(1+(xy)^2)

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

The principle is the same as before. There are just a few rules for finding the derivative: power rule, chain rule, product rule, quotient rule, etc. I will take just one term from your equation and differentiate: x^4y x^4y' + 4x^3y' For the last term: if arctan(A) = B, then it means A = tan(B)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

i am asking if u if the derivative of 5+x^4y=4x^3y+x^4y'? also if the derivative of arctan(xy)=1/(1+(xy)^2)

OpenStudy (ranga):

Oh, I thought that was the original equation they gave. Yes the derivative of 5+x^4y is 4x^3y+x^4y' Give me a minute for the other one.

OpenStudy (ranga):

derivative of arctan( something ) = 1 / { 1 + (something)^2 } times derivative of ( something )

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so derivative of arctan(xy)=(1/(1+(xy)^2))(y')

OpenStudy (ranga):

The 1 / { 1 + (xy)^2 } part is correct. But the derivative of xy: xy' + y

OpenStudy (anonymous):

ok

OpenStudy (ranga):

alright.

OpenStudy (anonymous):

i tried to figure it out myself but i cant arctan(xy)=5+x^4y

OpenStudy (anonymous):

i just need to find the derivative

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

You need derivative of arctan(xy) ?

OpenStudy (anonymous):

i need the derivative of the whole thing

OpenStudy (ranga):

Is the question: find dy/dx or y' ?

OpenStudy (anonymous):

arent they the same thing

OpenStudy (anonymous):

well its the same as before

OpenStudy (ranga):

They are the same thing. But I am asking what is the question? You can differentiate the right hand side and you can differentiate the right hand side. That will just leave some expression on the right = some expression on the right. But what are they asking you to find?

OpenStudy (anonymous):

y' of the whole thing

OpenStudy (ranga):

Okay that makes sense. Differentiate the left side. We just did the arctan thing earlier. Differentiate the right side. Group all y' terms and then express it as y' = some expression

OpenStudy (anonymous):

it doesnt specify whether we have to use implicit differentiation or not. but i got 1/(1+(xy)^2)(xy+y')=4x^3y+x^4y

OpenStudy (anonymous):

\[1/(1+(xy)^2)(xy'+y)=4x^3y+x^4y\]

OpenStudy (ranga):

The last term should be x^4y'

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and is it (xy+y') or (xy'+y)

OpenStudy (anonymous):

at the beginning

OpenStudy (ranga):

derivative of xy is: xy' + y (just product rule: first term times derivative of second term + second term times derivative of the first term)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would i do it

OpenStudy (ranga):

(xy' + y) / (1+ (xy)^2) = 4x^3y + x^4y' Multiply throughout by (1+ (xy)^2) Group y' terms Express it as y' = some expression

OpenStudy (anonymous):

it would be (1+x^2y^2)

OpenStudy (anonymous):

how would u group the y'

OpenStudy (ranga):

Instead of typing the whole thing which is time consuming I am going to call (1+ (xy)^2) as t (xy' + y) / t = 4x^3y + x^4y' Multiply throughout by t (xy' + y) = 4tx^3y + tx^4y' xy' - tx^4y' = 4t^3y - y = y(4t^3 - 1) y'(x - tx^4) = y(4t^3 - 1) y' = y(4t^3 - 1) / (x - tx^4) put t back as (1+ (xy)^2)

OpenStudy (anonymous):

y(4(1+(xy)2)3−1)x−(1+(xy)2)x4 this is what i plugged in and i got it wrong

OpenStudy (ranga):

If I expand all the terms and then factor it out I get:\[\Large y' = \frac{ y(-4x ^{5}y^{2} - 4x^{3} + 1) }{ x(x^{5}y^{2} + x^{3} - 1) }\] The method is correct. But if the answer doesn't agree it must be some trivial error and rechecking will catch the error.